OpenStudy (deadshot):

What is the simplified form of 2 over x squared minus x minus 1 over x ?

OpenStudy (deadshot):

\[\frac{ 2 }{ x^2 }-\frac{ x-1 }{ x }\]

OpenStudy (anonymous):

\[\frac{2x-(x-1)}{x^2}={x+1\over x^2}\]